OpenStudy (anonymous):

how do i subtract 4 and 3 7/8?

OpenStudy (gorv):

\[\huge 3 7/8=\huge 3+\frac{ 7 }{ 8 }\]

OpenStudy (gorv):

\[4-(3+\frac{ 7 }{ 8 })=4-3-\frac{ 7 }{ 8 }\]

OpenStudy (gorv):

can u do it now